The speed of a car increased from 36km/hr to 96km/hr in 0.6km . How long did this take and what was the acceleration

1
Expert's answer
2022-03-15T10:35:50-0400

The acceleration of the car

"a=\frac{v_f-v_i}{t}"

Hence, the time of motion

"t=\frac{v_f-v_i}{a}"

"t=\frac{96-36}{0.6}=100\:\rm hr"